OpenStudy (anonymous):

x^3 + 14x^2 - 32x = 0

OpenStudy (saifoo.khan):

\[-16, 0 ,2. \]

OpenStudy (anonymous):

thank you but could you help me with the steps

OpenStudy (anonymous):

start with \[x(x^2+14x-32)=0\] and the factor the second part if you can

OpenStudy (anonymous):

k

OpenStudy (anonymous):

\[x(x-2)(x+16)=0\] etc

OpenStudy (saifoo.khan):

factor them. x(x^2 +14 - 32) \[x(x-2) (x+16)\] x = 0 x = 2 x = -16

OpenStudy (anonymous):

what dexter said